# A First-look at Spatially-resolved Infrared Supernova remnants in M33 with JWST

### v1

@@ 3.6 Relation with progenitor masses of SNRs
In Figure 17, we show the median and 68th percentile progenitor masses of the SNRs based on the recent star-formation histories (56 Myr, or 7.1 M⊙) around the SNRs, as a proxy for the amount of young star formation in the regions. As a reminder from Section 2.4, these masses are from Koplitz et al. (2023) from the age-range of the local stellar population from the PHATTER survey, using stellar age-to-mass relations from single-stellar evolution models. We denote SNRs of particular interest in darker shades, such as the five clearly detected SNRs in multiple JWST filters (dark blue), and the MIRI partially-detected SNRs (light blue). The possible PAH-emitting L10-060 in the F770W-strip, and narrowband-filter detected L10-118 and L10-124 in the NGC 604 field are shown in grey. Both the grid-based and best-fit stellar masses are shown in the Figure.
@@
It is worth noting however, that all the clearly-detected SNRs – L10-039, 045, 071, 036 and 124 – have significant recent (56 Myr) star-formation (and thus progenitor masses 8 M⊙) in their vicinity. In contrast, some of the SNRs in the partial/non-detection categories such as L10-64, 78, 90, have best-fit masses 8 M⊙. In general, SNe from higher mass progenitors will be expected to go off in regions of denser ISM, but overall Figure 17 indicates that the infrared visibility of SNRs is not as strongly dependent on progenitor mass/age as it is on the local ISM conditions, age of the shock and/or presence of confusing background emission. This weak dependence on progenitor age is consistent with the observation that most SNRs in galaxies are in their Sedov/post-Sedov stages of evolution (Gordon et al., 1999; Badenes, 2010), where interaction with the swept-up ISM is primarily driving the shock dynamics and observed emission, instead of the initial ejecta.
Figure 17 shows that there is no dramatic difference in masses between the JWST-detected and non-detected/ambiguous-detection SNRs. In the MIRI field, for example, SNRs with high-mass progenitors (15 M⊙) can be bright and clearly detected with JWST like the cases of L10-071 and L10-124, as well as ambiguous like the cases of L10-081 and L10-083 (these two are particularly affected by bright background emission). There is also noticeable variation in progenitor properties within the clearly-detected SNRs. For example L10-039, which is the brightest SNR in the sample at all wavelengths, has a lower mass progenitor of M⊙ than L10-071 with best-fit mass of M⊙. The clearly-detected SNR with the highest mass progenitor is L10-124 with 35.5 M⊙.
@@ 4.1 Near and mid-infrared detection statistics
Visual inspection of the SNRs in the sharp JWST images revealed 3 clear and 6 partial detections out of 21 SNRs in the MIRI field, giving a clear detection fraction of 14% and total detection fraction of 43%. These numbers are roughly consistent with mid-IR detection statistics of SNRs elsewhere in the Local Group. Matsuura et al. (2022) found that out of the 24 known SNRs in the SMC, 5 have clear detections and 6 have possible detections in Spitzer IRAC and MIPS images, giving a clear detection fraction of 21% and total detection fraction of 46%. Similarly Galactic SNRs have shown detection rates of 20-30% with IR instruments, although the Galaxy has a greater degree of line-of-sight confusion from overlapping sources (e.g. Saken et al., 1992; Reach et al., 2006; Chawner et al., 2019). LMC SNRs show a higher detection fraction of about 62% (29 out of 47 SNRs detectable at 24 m) according to the work of Seok et al. (2013). This could be due to LMC being an external galaxy with lower line of sight confusion than the Galactic midplane. Physical differences in the environments (e.g. densities, metallicities and dust properties) of the LMC SNRs compared to SNRs in other Local Group galaxies could also be responsible for the different detection fractions (e.g. see discussions in Seok et al., 2013; Maggi et al., 2016; Auchettl et al., 2019).
@@
The IR emission in our clearly-detected SNRs is most likely dominated by forward shock interaction with the surrounding ISM. This is supported by the fact that their IR morphologies match the full extent of their optical emission-line structures (Figures 5, 8). Further evidence that these SNRs are past the ejecta-dominated phase comes from analyses of their X-ray spectra and sizes, which indicate swept-up masses 100-200 M⊙ (Long et al., 2010; Garofali et al., 2017), as well as the lack of any broad lines from high-velocity (1000 km/s) ejecta that has been observed in well-known ejecta-dominated SNRs such as Cas A, E0102, and N132D (Long et al., 2010, 2018),
@@
The clearly-detected JWST SNRs are likely interacting with denser ambient gas, as indicated by the [S II]-doublet ratios, and as a consequence are also among the smallest, and brightest in other wavelengths, compared to the rest of the sample. These characteristics are consistent with observations of many interacting SNRs in other galaxies. For example, N63A, N49 and N103B are among the brightest 24 m-emitting SNRs in the LMC (Seok et al., 2013), the smallest ones (Badenes, 2010), and have published optical spectroscopy that show [S II]:6717/6731 of 0.5-1 (Russell & Dopita, 1990; Li et al., 2021). The presence of dense ambient gas has also been independently confirmed by interaction signatures with molecular clouds (Oliva et al., 1989; Banas et al., 1997; Sano et al., 2018, 2019, 2023). Many of the IR-brightest MW SNRs also share these characteristics (e.g. Reach et al., 2006; Pinheiro Gonçalves et al., 2011), and also have independent evidence of interaction with dense nearby clouds (e.g. Jiang et al., 2010; Slane et al., 2015; Kilpatrick et al., 2016, and references therein). The mid-IR-spectra of interacting SNRs in the Galaxy like W28, W44, and 3C391 indicate that emission in the F560W filter (spanning 5-6 m) would most likely be dominated by prominent line emission of [Fe II] 5.34 m and H2 S(7) 5.51 m, while the emission in F2100W (18.7-23 m) is produced by continuum emission from collisionally-heated dust (e.g. Reach et al., 2002; Neufeld et al., 2007; Hewitt et al., 2009).
@@ 4.3 Evidence of shocks interacting with molecular gas
Our findings indicate that narrowband MIRI and NIRCam filters can be powerful tools for identifying SNRs interacting with cold, dense clouds in the ISM in external galaxies as far as M31 and M33. Both L10-118 and L10-124 SNRs provide the most direct evidence of this, with significant emission in the F470N filters that contain the H2 and CO lines (Figures 1113). Signatures of dense ISM traced by ALMA CO and F335M at the location of the molecular emission-rich shocks further corroborate the presence of interaction. These are also the farthest SNRs (to our knowledge) where such shocks with molecular emission have been found. Additionally, L10-124 shows evidence of the same molecular shock arc in F200W and F770W, which contain other well known ro/vibrational lines of H2 such as 1-0 S(1) 2.12 m, 0-0 S(5) 6.91 m, and 0-0 S(4) 8.03 m. Bright H2 lines have been observed in several SNRs interacting with molecular clouds (e.g. Reach & Rho, 2000; Andersen et al., 2011; Rho et al., 2021), and are typically produced in shocks 100 km/s traversing gas clouds with densities of 103 cm-3 (e.g. Hollenbach & McKee, 1979, 1989; Kaufman & Neufeld, 1996; Lehmann et al., 2020, 2022; Kristensen et al., 2023). Such molecule-rich shocks may also be present in the other clear and partially-detected SNRs, many of which are in close proximity to dense molecular clouds (Figure 3).
@@ 4.4 Effect of shocks on PAHs
We briefly comment here about PAH emission from SNRs, which is still a somewhat poorly-understood topic. Only five SNRs in our survey (the F770W-strip objects in Figure 15 and L10-124 in Figure 11) were observed in the F770W filter. We find a hint of a trend similar to the MIRI SNRs: The smaller SNRs L10-060 and L10-124 appear to show a somewhat brighter patch of emission within their optical-emitting region. L10-124 in particular also shows a clear bright F770W-emitting patch in Figure 11 coincident with its bright molecular shock that we identified in the NIRCam filters (hints of the 3.3 m PAH feature are also visible in F335M, Figure 9). While we concluded in Section 3.3 that this could be from additional H2 rotational line series in the F770W bandwidth, since the shape of the F770W arc is almost the same as in the NIRCam images, some contribution from swept-up PAH-rich gas or pre-shock PAH gas irradiated by the precursor could still exist, as has been seen from spectroscopy and imaging of some young, interacting SNRs (Tappe et al., 2006; Neufeld et al., 2007; Tappe et al., 2012).
@@ 5 Conclusions
In this paper, we present the first JWST views of a subset of SNRs in M33 showcasing the detailed near/mid-IR structures of extragalactic SNRs that can be revealed in the sub-arcsecond resolution images. We pulled observations from four regions of M33 that were observed with NIRCam and MIRI filters during Cycles 1 and 2. These are: 1) “MIRI field” – a kpc2 mosaic of the southern arm of M33, observed with F560W (5.6 m) and F2100 (21 m) filters. 2) “NIRCam field” – a kpc2 mosaic of the central-eastern part of M33, observed with F335M (3.35 m) and F444W (4.4 m) filters. 3) “F770W-strip” – a roughly 5 kpc long strip outwards from the center of M33, observed with the F770W filter, and 4) “NGC 604” – mosaic of the giant HII region NGC 604 observed with several NIRCam and MIRI broadband and narrowband filters. A combined 43 SNRs fall in these fields, spanning a variety of sizes and ISM environments.
We visually inspect whether these SNRs are detectable in the various JWST filters, using their morphologies in narrow-band optical images from LGGS as a reference, and compare the various categories of detection with the known multi-wavelength properties of the SNRs. Our analysis reveals the following:
@@
We estimate, by eye, that about 43% of the SNRs in the MIRI field have complete or partial detections of their optical morphologies in the MIRI filters. NIRCam only has 1 clearly detected SNR out of 15, so a detection fraction of only 6.7%. These detection statistics are similar to IR SNRs in the SMC, smaller than LMC, but higher than the Milky Way. Considering that M33 is farther than all three, our results further endorse JWST’s exquisite sensitivity and resolution for surveying extragalactic SNRs beyond the Magellanic Clouds, marking a major leap over the capabilities of the previous-generation Spitzer (Figure 1).
@@
Three SNRs in the MIRI field – L10-039, L10-045 and L10-071 – have infrared morphologies that span the full extent of their optical morphologies (Figure 5), signifying strong contribution from emission at their forward shocks in the infrared filters. They are among the brightest SNRs in M33, and are prominent even in broadband HST images, covering many of the expected forbidden lines in the optical (e.g. H, [O III]). Six SNRs in the MIRI field have partial detections (e.g. a single filament in L10-069, part of the shell in L10-058 etc., see Figure 6), while the rest have no detectable emission or emission that is unique from the background (Figure 7).
@@
The NIRCam field only has 1 clearly-detected SNR – L10-036 – clearly visible in F335M and F44W. The prominent ear-shaped morphology extends to the near-IR as well, and is likely shaped by a dense, asymmetrical environment as evidenced from ALMA images. The other 14 SNRs do not have any detectable emission in the filters (though it is worth noting these are not yet stellar-subtracted). This is consistent with the lower detectability of SNRs at shorter IR wavelengths seen in the Magellanic Clouds and Milky Way.
@@
Three SNRs in NGC 604 have multiple NIRCam and MIRI filter images (Figure 10). SNRs L10-118 and L10-124 show significant emission in F470N, signifying excitation of H2 (and possibly some CO) due to interaction with molecular gas visible in ALMA (Figures 11, 13). These are the farthest SNRs where such molecule-rich shocks have been discovered. The smallest SNR L10-124 is visible across 0.9-7.7m, although with diminishing visibility in MIRI (F1130W, F1500W). The largest SNR LL14-168 (80 pc) shows faint emission only in the F187N filter (Pa), and evidence of swept-up ambient ISM in F335M and ALMA CO (Figure 14).
@@
Five SNRs in our sample were observed in F770W, giving insight into the effect of shocks on the PAH-rich ISM. Smaller SNRs like L10-060 and L10-124 show patches of elevated emission in these filters possibly due to heating of the PAH, while the larger SNRs like L10-65 have more depleted regions of PAH within them. The prominent molecular shock arc of L10-124 is visible in F770W, likely captured by the S(5) and S(4) transition lines, but they are not visible in F1130W and F1500W.
@@
We do not find any strong correlation between visibility of the SNRs in JWST and progenitor masses inferred by Koplitz et al. (2023) from the local star-formation histories from PHATTER survey. Among the prominent JWST SNRs, the highest mass progenitor is L10-124 (35 M⊙) and the lowest mass is L10-036 (9.7 M⊙). This reinforces the idea in the previous point that infrared visibility is more strongly affected by the local ISM conditions, age of the shock, and overlapping background emission.
@@ 6 Acknowledgements
Appendix A Individual descriptions of SNRs in the MIRI field
@@
L10-039 – Brightest IR SNR in the MIRI field, 2nd brightest H SNR. In the JWST images (and also in HST), it appears to be a system of an outer and a high-surface brightness inner shock of about 0.63″( 2.57 pc) diameter. Could be shock interaction with an inner ring of dense material. The morphology is almost spherically symmetric but incomplete, particularly in the northwest region where the rims show discontinuity. Similar loops seen in W49B. 21 m however fills the whole region. Larger arcs present as noted by Long et al. (2010) though the possibility of overlapping SNR cannot be ruled out. Located on the edge of a large dust-cleared bubble of 108185 pc diameter.

### v2

@@ 3.6 Progenitor masses of Southern Arm SNRs
Given the distinct range of densities and velocities spanned by the clearly-detected SNRs compared to the other SNRs in the Southern Arm (Section 3.5 and Figure 16), it would also be interesting to check if they share a similar distinction in progenitor properties. We check this in Figure 17 using the median and 68th percentile progenitor masses of the SNRs in the Southern Arm based on the recent star-formation histories (56 Myr, or 7.1 M⊙) around the SNRs, as a proxy for the amount of young star formation in the regions. As a reminder from Section 2.4, these masses were measured by Koplitz et al. (2023) from the age-range of the local stellar population from the PHATTER survey, using stellar age-to-mass relations from single-stellar evolution models. We show both grid-based masses, which are measured from the SFH map of Lazzarini et al. (2022) at the location of the SNR, and best-fit masses, which were measured by Koplitz et al. (2023) in the 50 pc SNR vicinity, after removing the background SFH (see Section 2.4 for more details). We denote SNRs of particular interest, i.e. the clearly-detected and partially-detected populations, in two shades of blue.
Figure 17 shows that there is no conspicuous difference in progenitor mass range of the three categories of SNRs in the Southern Arm. For example, SNRs with high-mass progenitors (15 M⊙) can be bright and clearly detected with JWST like the cases of L10-071, as well as ambiguous like the cases of L10-081 and L10-083 (these two are particularly affected by bright background emission). There is also noticeable contrast in progenitor properties within the clearly-detected SNRs. For example L10-039, which is the brightest SNR in the sample at all wavelengths, has a lower mass progenitor of M⊙ than L10-071 with best-fit mass of M⊙. It is worth noting however, that all the clearly-detected SNRs – L10-039, 045, 071 – have significant recent (56 Myr) star-formation (and thus progenitor masses 8 M⊙) in their vicinity. In contrast, of the 18 partially-detected and ambiguous SNRs, 5 have best-fit masses 8 M⊙. These SNRs were described by Koplitz et al. (2023) as possible but ‘weaker’ Type Ia candidates, because their grid-based masses in Figure 17 still indicates the presence of 56 Myr star-formation nearby, just not in their immediate 50 pc vicinity. These SNRs, however, could also have been produced by a core-collapse progenitor that originated in a star-formation site farther than 50 pc, and drifted to the current SNR location before exploding. The progenitor could also be from the delayed (50-200 Myr) core-collapse channel predicted by Zapartas et al. (2017) from mergers of low mass (8 M⊙) progenitors.
@@
We can thus conclude that whether the SNRs are clearly or even partially-detected in infrared is mostly independent of the progenitor, and more directly correlated with the ISM conditions and age of the shock (as shown in Figure 16). This is generally consistent with the observation that most SNRs in galaxies are in their Sedov/post-Sedov stages of evolution (Gordon et al., 1999; Badenes, 2010), where interaction with the swept-up ISM is primarily driving the shock dynamics and observed emission, instead of the initial ejecta.
@@ 4.1 Near and mid-infrared detection statistics
Visual inspection of the SNRs in the sharp JWST images revealed 3 clear and 6 partial detections out of 21 SNRs in the Southern Arm field, giving a clear detection fraction of 14% and total detection fraction of 43%. These numbers are roughly consistent with mid-IR detection statistics of SNRs elsewhere in the Local Group. Matsuura et al. (2022) found that out of the 24 known SNRs in the SMC, 5 have clear detections and 6 have possible detections in Spitzer IRAC and MIPS images, giving a clear detection fraction of 21% and total detection fraction of 46%. Similarly Galactic SNRs have shown detection rates of 20-30% with IR instruments, although the Galaxy has a greater degree of line-of-sight confusion from overlapping sources (e.g. Saken et al., 1992; Reach et al., 2006; Chawner et al., 2019). LMC SNRs show a higher detection fraction of about 62% (29 out of 47 SNRs detectable at 24 m) according to the work of Seok et al. (2013). This could be due to LMC being an external galaxy with lower line of sight confusion than the Galactic midplane. Physical differences in the environments (e.g. densities, metallicities and dust properties) of the LMC SNRs compared to SNRs in other Local Group galaxies could also be responsible for the different detection fractions (e.g. see discussions in Seok et al., 2013; Maggi et al., 2016; Auchettl et al., 2019).
4.2 What is driving broadband emission in the clearly-detected SNRs?
@@
The IR emission in our clearly-detected SNRs is most likely dominated by forward shock interaction with the surrounding ISM. This is supported by the fact that their IR morphologies match the full extent of their optical emission-line structures (Figures 5, 8). The SNRs are likely in their Sedov phases given their X-ray spectra and sizes, which indicate swept-up masses 100-200 M⊙ (Long et al., 2010; Garofali et al., 2017), and their optical spectra that lack signatures of high-velocity (1000 km/s) ejecta that has been observed in well-known ejecta-dominated SNRs such as Cas A, E0102, and N132D (Long et al., 2010, 2018).
@@
The presence of dense ambient media around the clearly-detected SNRs is characteristic of other well-known infrared-bright interacting SNRs. For example, N63A, N49 and N103B are among the brightest 24 m-emitting SNRs in the LMC (Seok et al., 2013), the smallest ones (Badenes, 2010), and have published optical spectroscopy that show [S II]:6717/6731 of 0.5-1 (Russell & Dopita, 1990; Li et al., 2021). The presence of dense ambient gas has also been independently confirmed by interaction signatures with molecular clouds (Oliva et al., 1989; Banas et al., 1997; Sano et al., 2018, 2019, 2023). Many of the IR-brightest MW SNRs also share these characteristics (e.g. Reach et al., 2006; Pinheiro Gonçalves et al., 2011), and also have independent evidence of interaction with dense nearby clouds (e.g. Jiang et al., 2010; Slane et al., 2015; Kilpatrick et al., 2016, and references therein). Spectroscopic observations of many of these local SNRs suggest that the F560W emission of the clearly-detected SNRs is likely driven by prominent lines of [Fe II] 5.34 m and H2 S(7) 5.51 m, while the emission in F2100W (18.7-23 m) is produced by continuum emission from collisionally-heated dust (e.g. Reach et al., 2002; Neufeld et al., 2007; Hewitt et al., 2009).
@@ 4.4 Evidence of shocks interacting with molecular gas
Our findings indicate that narrowband MIRI and NIRCam filters can be powerful tools for identifying SNRs interacting with cold, dense clouds in the ISM in external galaxies as far as M31 and M33. Both SNRs L10-118 and L10-124, which are in the NGC-604 field, provide the most direct evidence of this, with significant emission in the F470N filters that contain the H2 and CO lines (Figures 1113). Signatures of dense ISM traced by ALMA CO and F335M at the location of the molecular emission-rich shocks further corroborate the presence of interaction. These are also the farthest SNRs (to our knowledge) where such shocks with molecular emission have been found. Additionally, L10-124 shows evidence of the same molecular shock arc in F200W and F770W, which contain other well known ro/vibrational lines of H2 such as 1-0 S(1) 2.12 m, 0-0 S(5) 6.91 m, and 0-0 S(4) 8.03 m. Bright H2 lines have been observed in several SNRs interacting with molecular clouds (e.g. Reach & Rho, 2000; Andersen et al., 2011; Rho et al., 2021), and are typically produced in shocks 100 km/s traversing gas clouds with densities of 103 cm-3 (e.g. Hollenbach & McKee, 1979, 1989; Kaufman & Neufeld, 1996; Lehmann et al., 2020, 2022; Kristensen et al., 2023). Such molecule-rich shocks may also be present in the other clear and partially-detected SNRs, many of which are in close proximity to dense molecular clouds (Figure 3).
@@ 4.5 Effect of shocks on PAHs
We briefly comment here about PAH emission from SNRs, which is still a somewhat poorly-understood topic. Only five SNRs in our survey (the North Radial Strip objects in Figure 15 and L10-124 in Figure 11) were observed in the F770W filter. We find a hint of a trend similar to the MIRI SNRs: The smaller SNRs L10-060 and L10-124 appear to show a somewhat brighter patch of emission within their optical-emitting region. L10-124 in particular also shows a clear bright F770W-emitting patch in Figure 11 coincident with its bright molecular shock that we identified in the NIRCam filters (hints of the 3.3 m PAH feature are also visible in F335M, Figure 9). While we concluded in Section 3.3 that this could be from additional H2 rotational line series in the F770W bandwidth, since the shape of the F770W arc is almost the same as in the NIRCam images, some contribution from swept-up PAH-rich gas or pre-shock PAH gas irradiated by the precursor could still exist, as has been seen from spectroscopy and imaging of some young, interacting SNRs (Tappe et al., 2006; Neufeld et al., 2007; Tappe et al., 2012).
@@ 5 Conclusions
In this paper, we present the first JWST views of a subset of SNRs in M33 showcasing the detailed near/mid-IR structures of extragalactic SNRs that can be revealed in the sub-arcsecond resolution images. We pulled observations from four regions of M33 that were targeted with NIRCam and MIRI filters during Cycles 1 and 2. A combined 40 SNRs fall in these fields, spanning a variety of sizes and ISM environments. We visually inspect whether these SNRs have complete or even partial detections of their optical morphologies (from LGGS survey) in JWST, and whether there is any correlation between their infrared detections and published multi-wavelength properties. Our analysis reveals the following:
@@
We estimate, by eye, that about 43% of the SNRs in the Southern Arm field have complete or partial detections of their optical morphologies in the MIRI filters, and only 6.7% (1 out of 15 detections in the Center field) in NIRCam filters. These detection statistics are similar to near/mid-IR observations of SNRs in the SMC, but midway between those in the LMC and Milky Way.
@@
Three SNRs in the Southern Arm field – L10-039, L10-045 and L10-071 – have infrared morphologies nearly identical to their optical morphologies (Figure 5). They are also prominent in broadband HST filters that capture many expected forbidden lines in the optical (e.g. H, [O III]). Six SNRs in the Southern Arm field have partial detections (e.g. a single filament in L10-069, part of the shell in L10-058 etc., see Figure 6), while the rest have no detectable emission or emission that is unique from the background (Figure 7).
@@
The Center field only has 1 clearly-detected SNR – L10-036 – clearly visible in F335M and F44W. The prominent ear-shaped morphology extends to the near-IR as well, and is likely shaped by a dense, asymmetrical environment as evidenced from ALMA images.
@@
Of the three SNRs in NGC 604 with images in multiple NIRCam and MIRI filters (Figure 10), L10-118 and L10-124 show significant emission in F470N, signifying excitation of H2 (and possibly some CO) due to interaction with molecular gas visible in ALMA (Figures 11, 13). These are the farthest SNRs where such molecule-rich shocks have been discovered. The smallest SNR L10-124 is visible across 0.9-7.7m, although with diminishing visibility in MIRI (F1130W, F1500W). The largest SNR LL14-168 (80 pc) shows faint emission only in the F187N filter (Pa), and evidence of swept-up ambient ISM in F335M and ALMA CO (Figure 14).
@@
Of the five SNRs observed with F770W (tracing PAH) in our survey, the smaller SNRs like L10-060 and L10-124 show patches of elevated emission, while the larger SNRs like L10-65 appear to have emission voids within them. The prominent molecular shock arc of L10-124 is visible in F770W, likely captured by the S(5) and S(4) transition lines, but they are not visible in F1130W and F1500W.
@@
We do not find a strong correlation between the local (50 pc) star-formation history-based progenitor masses of SNRs from Koplitz et al. (2023) and whether they are clearly or partially-detected or undetected in JWST (Section 3.6). All three categories of SNRs plausibly span the progenitor mass range of 8 and 30 M⊙, based either on local star-formation histories measured by Koplitz et al. (2023), or from the M33 star-formation history map of Lazzarini et al. (2022).
@@ 6 Acknowledgements
Appendix A Individual descriptions of SNRs in the Southern Arm field
@@
L10-039 – Brightest IR SNR in the Southern Arm field, 2nd brightest H SNR. In the JWST images (and also in HST), it appears to be a system of an outer and a high-surface brightness inner shock of about 0.63″( 2.57 pc) diameter. Could be shock interaction with an inner ring of dense material. The morphology is almost spherically symmetric but incomplete, particularly in the northwest region where the rims show discontinuity. Similar loops seen in W49B. 21 m however fills the whole region. Larger arcs present as noted by Long et al. (2010) though the possibility of overlapping SNR cannot be ruled out. Located on the edge of a large dust-cleared bubble of 108185 pc diameter.